OpenStudy (anonymous):

What is the value of the expression |x + y| when x = -7 and y = 18? A. -25 B. -11 C. 11 D. 25 What is the mean of the set of numbers? -10, -5, 2, 6, 17 What is 451.538 rounded to the tenths place? Which of the following lists all the integer solutions of the inequality |x| < 5? 4, -4 0, 1, 2, 3, 4 - 4, -3, -2, -1, 0, 1, 2, 3, 4 - 5, -4, -3, -2, -1, 0, 1, 2, 3, 4, 5

OpenStudy (anonymous):

|x + y| --- x = -7 and y = 18 | -7 + 18| | 11 | answer is 11 because absolute values are positive the mean is the average, so we must add them... -10 + (-5) + 2 + 6 + 17 = -10 - 5 + 2 + 6 + 17 = -15 + 25 = 10 then we divide by 5 (because there are 5 numbers) 10/5 = 2 (the mean) 451.438 = 451.4 the solution would be x < 5 and x > -5 answer is : last answer choice

OpenStudy (anonymous):

|x+y| plug in |-7+18| Whatever value you get make it positive mean = sum of all values / sum of number of values \[\frac{ -10 + -5 + 2 + 6 + 17 }{ 5 }\] 451.538 5 is the tenths place So look one to the right of 5 and determine if that is {0...4} to keep as 5 or {5...9} to round up to 6. |a| < b splits into a < b and a > -b which combine into -b < a < b Do it with your problem. Fill in all the integers that work.

OpenStudy (the_fizicx99):

@kelliegirl33 wouldn't the last one be c because 5 isn't < 5 ? :o

OpenStudy (anonymous):

your right...I missed that. It is C because there is no equal sign in the problem to include -5 and 5. I am sorry about that. Good catch :)

OpenStudy (anonymous):

OK thank you so much!!!

OpenStudy (anonymous):

YW :)

OpenStudy (the_fizicx99):

Np :}